OpenStudy (anonymous):

If y = 9x − 7, which of the following sets represents possible inputs and outputs of the function, represented as ordered pairs? A. {(0,-7),(1,2),(-1,-16)} B. {(-7,0),(2,1),(-16,-1)} C. {(1,9), (2,7), (3,16)} D. {(7,9), (8,10), (9,11) HELLPP!!!

OpenStudy (anonymous):

@Nnesha

Mehek (mehek14):

substitute the values

OpenStudy (anonymous):

how

Mehek (mehek14):

it would be the first one because -7= 9*0 - 7 is 0 2 = 9*1 -7 IS 9 -16 = 9*-1 -7 is -9

Mehek (mehek14):

substitute the x values for x and the y values for y

OpenStudy (anonymous):

thank you so much

Mehek (mehek14):

no problem